# Learning the Geometric Mechanics of Robot Motion Using Gaussian Mixtures

### v1

@@ Abstract
Data-driven models of robot motion constructed using principles from Geometric Mechanics have been shown [[bittner2018geometrically, zhao2022walking, hatton2013geometric]] to produce useful predictions of robot motion for a variety of robots. For robots with a useful number of DoF, these geometric mechanics models can only be constructed in the neighborhood of a gait. Here we show how Gaussian Mixture Models (GMM) can be used as a form of manifold learning that learns the structure of the Geometric Mechanics “motility map [^1] In the geometric mechanics literature, what we refer to as the “motility map” here is often called the “(negative of the) local connection”. This terminology reflects its mathematical provenance from the fiber bundle literature, and not its application. We have therefore chosen a more readable name for use here.” and demonstrate: [i] a sizable improvement in prediction quality when compared to the previously published methods; [ii] a method that can be applied to any motion dataset and not only periodic gait data; [iii] a way to pre-process the data-set to facilitate extrapolation in places where the motility map is known to be linear. Our results can be applied anywhere a data-driven geometric motion model might be useful.
I Introduction
I-A A Brief Review of Classical Geometric Mechanics
Consider a collection of inter-connected bodies whose configuration $q$ is subject to forces $F(q,\dot{q})$ that are equivariant under changes of position and orientation due to a Lie group $\mathfrak{G}$, such as $\mathfrak{G}=\mathsf{SE}(3)$. The equivariance means that $F(g(q),ad_{g}(\dot{q}))=ad^{*}_{g}F(q,\dot{q})$ i.e. for any group element $g\in\mathfrak{G}$, i.e. changes in coordinate frame transform forces to the forces that would have arisen from the transformed positions and velocities. Typically this arises by having multiple body parts whose positions relative to the body frame are defined by its “shape”, and having the forces arise through contact interactions of those body parts with a homogeneous environment. From a mathematical perspective, it is “shape” that arises from the equivalence of body configurations under the Lie group action, i.e. shape is the equivalence class $r:=[q]_{\mathfrak{G}}$. The configuration thus partitions into a changing “(body) shape” $r$ and a moving body frame $g$, with associated body velocity in the relevant Lie algebra, $v_{b}\in\mathcal{A}_{\mathfrak{G}}$ (e.g. $\mathsf{se}(3)$ for $\mathsf{SE}(3)$).
@@ I-A A Brief Review of Classical Geometric Mechanics
for some momentum $p$ associated with the moving body frame, and expressed through a shape dependent “frozen inertia tensor” $\mathbb{I}(r)$. For many systems, $p=0$, or $p\to 0$ so quickly that the $p=0$ assumption is reasonable, leaving only $A(r)$ to govern the motion; such systems are called “principally kinematic”. This $A(r)$ term is known as the “local connection” or “motility map”.
In previous work our group has shown methods that can learn the motility map model in the neighborhood of a periodic motion [[bittner2018geometrically]], and extended these learning methods to systems with non-zero momentum [[kvalheim2019gait]], and under-actuated systems with passive degrees of freedom [[bittner2022data]]. We have also shown that data driven motility map models are highly predictive for multilegged locomotion with or without slipping in both animals and robots [[zhao2022walking]]. Follow up work by Bittner also showed that these models could be updated with online learning [[deng2024adaptive]].
However, data driven motility map models remained either “easy” to identify near a periodic gait, or several related gaits [[bittner2021optimizing]], or “exponentially hard” to identify, requiring an amount of data exponential in the number of DoF for an exhaustive model [[dai2016geometric]] throughout the phase space.
I-B Contribution
In a principally kinematic system, the entire physics of body-environment interactions is captured by the motility map model $A(r)$ of the equation: $v_{b}=A(r)\dot{r}$. It is a mathematical truism that any function can be represented by its “graph” – the set of points that comprise its input-output pairing in the cartesian product of the domain and co-domain – and that graph is a manifold for smooth functions. The graph of the motility map can be represented by the constraint equation that is its indicator:
$f(v_{b},r,\dot{r})=v_{b}-A(r)\dot{r}=0.$ (2)
In general, one should exercise caution when learning manifolds, since without additional prior knowledge the local structure of a manifold at a point provides no information (other than dimension) about the local structure of the manifold at other points. Thus when learning a manifold it is typical to interpolate between known points of the manifold, but be cautious and conservative when extrapolating. However, the graph of a motility map is a “ruled surface” – it is a union of affine subspaces, because of its linearity in $\dot{r}$. This special structure suggests that at least these affine subspaces can be learned faster, e.g. by the use of an additional meta-parameter governing how much linear extrapolation can be allowed in the $\dot{r}$ direction(s).
@@ I-B Contribution
    We reformulate the problem of learning a motility map as a manifold learning problem, allowing for non-linearities in $\dot{r}$ such as those predicted by [[kvalheim2019gait]], other non-linear structures in $r$ and $\dot{r}$, and hysteretic multi-valued motility maps to be represented.
@@
    We provide a data transformation that allows the degree of linear extrapolation in the $\dot{r}$ direction to be controlled.
@@
    We present an adaptation of well-known Gaussian Mixture Model fitting techniques to this machine learning problem.
@@
    We demonstrate that when applied to the datasets of [[zhao2020multi]] (available at [[bigant6dataset]]), consisting of hexapedal robot data, this new approach noticeably improves the predictive ability relative to the prior state of the art.
II Methods
II-A Manifold Learning
Manifold learning is the process of estimating the structure of a manifold from a set of noisy observations taken from that manifold. It operates on the assumption that high-dimensional data often lies on or near a lower-dimensional manifold embedded within the higher-dimensional space. Since every smooth function defines a manifold which is its graph, methods of estimating manifolds from the input-output pairs of a function are themselves methods of function regession.
@@ II-A Manifold Learning
From a practical standpoint, many functions modeling physical systems exhibit hysteretic behavior – they can follow different branches of values for the same input, but once on a branch, they remain on that branch. This form of hysteresis can be represented computationally by looking up the value of a function based on both its parameters and the most recent value it returned, with the latter value serving to select the appropriate branch.
Here we suggest that the motility map’s graph be considered a manifold for such a manifold learning problem, as described by eqn.2.
II-B Gaussian Mixture Models as a form of Manifold Learning
This section describes why GMM modeling of sample distribution of the graph of a nonlinear function is a natural extension of “Total Least Squares” regression to manifolds.
Consider now a linear model $y=Ax$, $x\in X$, and $y\in Y$, applied to a sample of $x$ points coming from a Gaussian distribution $N(\mu,\sigma)$. The graph of this linear model is a point set in the direct sum vector space $W:=X\oplus Y$.
@@ II-B Gaussian Mixture Models as a form of Manifold Learning
We use the direct sum operator $\oplus$ to remind the reader that the vector space structure of the spaces $X$ and $Y$ extends naturally to $W$ as $(x\oplus y)+\alpha(v\oplus u):=(x+\alpha v)\oplus(y+\alpha u)$. Additionally, any linear map $M$ from $W$ to itself, i.e. $M\in\text{Aut}(W)$ has four parts: $M_{xx}\in\text{Aut}(Y)$, $M_{yy}\in\text{Aut}(Y)$, $M_{xy}\in L(Y,X)$ and $M_{yx}\in L(X,Y)$, such that $M(x\oplus y):=(M_{xx}x+M_{xy}y)\oplus(M_{yx}x+M_{yy}y)$ following the usual block structure of matrix operations. When $M_{xy}$ and $M_{yx}$ are both zero, we write the block diagonal $M$ as $M_{xx}\oplus M_{yy}$, noting that in this case $(M_{xx}\oplus M_{yy})(x\oplus y)=(M_{xx}x)\oplus(M_{yy}y)$.
@@
Thus the graph of points $(x,Ax)$ is itself a sample from a (degenerate) Gaussian $N(\mu\oplus A\mu,S)$, where $S$ is a rank $\dim X$ symmetric non-negative matrix whose $S_{xx}$ block is $\sigma$ and whose $S_{yy}$ block is $A\sigma A^{T}$. Assuming the measurements of $x$ (resp. $y$) are corrupted by (independent) Gaussian measurement noise with covariance $C_{x}$ (resp. $C_{y}$) the noisy data of the graph is still a sample from a Gaussian, now $N(\mu\oplus A\mu,~{}S+(C_{x}\oplus C_{y}))$, and no longer degenerate.
@@
Assuming $\sigma$ is sufficiently larger than $C_{x}$ and $C_{y}$, the matrix $A$ can be recovered from $S$ by using Total Least Squares regression. The TLS computation splits $W$ space along the eigenspaces of $S$, by decomposing $S=U^{T}dU$ with $U$ orthogonal and $d$ a non-negative diagonal matrix with non-increasing elements along the diagonal. The TLS estimate of the graph is the kernel of the $\dim Y$ last rows of $U$, i.e. the space $U_{yx}x+U_{yy}y=0$, or
@@
II-B1 Example
@@
Consider the algebraic variety $y^{2}=(x-1)^{2}(x+1)^{2}$ shown in Fig. 1. It consists of two branches: $y=(x-1)(x+1)$ and $y=-(x-1)(x+1)$, each of which is a manifold, and provides us with a toy example of a multi-valued function. The variety itself, as a point set in the $(x,y)$ plane, is of measure 0, but a sampling process sampling $x$ uniformly in $[-1.5,1.5]$, then taking a corresponding $y$ from either branch at equal probability, and finally corrupting both $x$ and $y$ with $N(0,0.15)$ Gaussian measurement noise, We built a GMM modeling these data, described in terms of a list of Gaussians $N(\mu_{k},S_{k})$ with their respective weights $w_{k}$.
Our predicted value for the function at $x$ with prior measurement $y_{p}$ is computed using as following. Define
$p_{k}(x,y)$ $:=p_{k}:=\Pr((x,y)|N(\mu_{k},S_{k})),$ (4)
the probability of the point $(x,y)$ belonging to the $k$-th Gaussian in the mixture model.
The predicted value $\hat{y}(x)$ given the previous value of $y_{p}$ is:
$\hat{y}(x):=\frac{\sum_{k}w_{k}p_{k}(x,y_{p})\left(A(x-\mu_{x,k})+\mu_{y,k}\right)}{\sum_{k}w_{k}p_{k}(x,y_{p})}$ (5)
@@
Figure 1: Example of Gaussian Mixture Model (GMM) based Manifold Regression. We considered the algebraic variety $y^{2}=(x-1)^{2}(x+1)^{2}$, which has symmetric positive (blue line) and negative (red line) branches. We simulated 1000 noisy measurements (blue circles) with measurement noise $N(0,0.15)$ in both $x$ and $y$ with $x$ taken uniformly from $[-1.5,1.5]$, and the $y$ branch chosen at random with equal probability. Using scikit.learn.gmm, we represented the sample points as a GMM and plotted the covariance ellipses of the constituent Gaussians (black ellipses). We then computed and plotted the predicted value of the model for a point traveling from $(-1,1)$ to the right, thus taking the positive (purple) branch, and similarly for a point traveling to the right from $(-1,-1)$ on the negative (dark red) branch.
II-C Pre-processing to Produce a Ruled Surface
As noted above, the GMM modeling approach makes no a-priori assumptions about the structure of the manifold being modeled, but the graph of the motility map is a ruled surface which is affine in $\dot{r}$.

### v2

# Learning Geometric Mechanics with A-GBR

@@ Abstract
Data-driven models of robot motion constructed using principles from Geometric Mechanics have been shown to produce useful predictions of robot motion for a variety of robots. For robots with a useful number of DoF, these geometric mechanics models can only be constructed in the neighborhood of a gait. Here we show how Gaussian Mixture Regression (GMR) can be used as a form of manifold learning that learns the structure of the Geometric Mechanics “motility map” [^2] In the geometric mechanics literature, what we refer to as the “motility map” here is sometimes called the “(negative of the) local connection”. This terminology reflects its mathematical provenance from the fiber bundle literature, and not its application. We have therefore chosen a more readable name for use here. and demonstrate: [i] a sizable improvement in prediction quality when compared to the previously published methods; [ii] a method that can be applied to any motion dataset and not only periodic gait data; [iii] a way to pre-process the dataset to facilitate extrapolation in places where the motility map is known to be linear. Our results can be applied anywhere a data-driven geometric motion model might be useful.
1 Introduction
Geometric Mechanics describes the decomposition of (Lagrangian) dynamics under a symmetry into “body frame” and “shape” coordinates. In environments where friction or viscosity annihilates body momentum sufficiently quickly, the consequence is a first-order equation of motion which is on its face quite unlike Newtonian dynamics – it is “geometric” in the sense that only the shape of the motion matters, but the speed does not.
1.1 A Brief Review of Classical Geometric Mechanics
Consider a collection of interconnected bodies whose configuration $q$ is subject to forces $F(q,\dot{q})$ that are equivariant under changes of position and orientation due to a Lie group $\mathfrak{G}$, such as $\mathfrak{G}=\mathsf{SE}(3)$. The equivariance means that $F(g(q),ad_{g}(\dot{q}))=ad^{*}_{g}F(q,\dot{q})$ i.e. for any group element $g\in\mathfrak{G}$, changes of the coordinate frame transform forces to the forces that would have arisen from the transformed positions and velocities. Typically this arises by having multiple body parts whose positions relative to the body frame are defined by its “shape”, and having the forces arise through contact interactions of those body parts with an environment which is homogeneous (outside perhaps the influences of the body on its immediate environs). From the mathematical perspective, “shape” is not the fundamental concept. Rather, it is “shape” that arises from the equivalence of body configurations under the Lie group action, i.e. shape is the equivalence class $r:=[q]_{\mathfrak{G}}$. The configuration $q$ thus partitions into a changing “(body) shape” $r$ and a moving body frame $g$, with an associated body velocity in the relevant Lie algebra, $v_{b}\in\mathcal{A}_{\mathfrak{G}}$ (e.g. $\mathsf{se}(3)$ for $\mathsf{SE}(3)$).
@@ 1.1 A Brief Review of Classical Geometric Mechanics
for some momentum $p$ associated with the moving body frame, and expressed through a shape dependent “frozen inertia tensor” $\mathbb{I}(r)$. For many systems, $p=0$, or $p\to 0$ so quickly that the $p=0$ assumption is reasonable, leaving only $A(r)$ to govern the motion; such systems are called “principally kinematic”, and the absence of $p$ from these equations lends “geometric mechanics” its name as “geometric”. The $A(r)$ term is known as the “local connection” or “motility map”, and estimating it from data is the subject of this paper.
1.2 Background
In 1989 Shapere and Wilczek published a physics paper that arguably started the field of geometric mechanics, by showing that shape-change in micro-organisms produced motion that is startlingly different from Newtonian mechanics intuitions (Shapere and Wilczek (1989)). In the time since, the locomotion of many biological and robotic systems across diverse terrains has been shown to be principally kinematic. Examples include locomotion on granular substrates like sand (Astley et al. (2020)), and multi-legged locomotion on solid ground (Zhao et al. (2022)). Morphologies studied include centipede-like robots, as in Zhao et al. (2022); Chong et al. (2023) and snake robots, as in Vaquero et al. (2024).
The complexity of body-environment interactions is hard to model in multi-contact or distributed contact situations. According to Astley et al. (2020), conventional approaches to multi-legged locomotion would require a combinatorial number of domains to model all contact states, and models of motion on sand would rely on having dynamical models for granular fluids which (so far) do not exist except in rare cases. While some groups are trying direct modeling approaches (Prasad et al. (2023); Wu et al. (2024)), data-driven methods are increasingly employed for modeling and control of principally kinematic systems (Deng et al. (2024); Rieser et al. (2024)).
Data-driven approaches are effective because the motility map is much easier to learn from data than more general cases of Newtonian dynamics. Once it is learned, various downstream tasks, such as control and gait optimization, can be efficiently performed as shown in Hatton and Choset (2011); Ramasamy and Hatton (2019); Yang et al. (2024).
In previous work, our group has demonstrated the means to learn the motility map model in the neighborhood of a periodic motion (Bittner et al. (2018)), and extended these learning methods to systems with non-zero momentum (Kvalheim et al. (2019)), and under-actuated systems with passive degrees of freedom (Bittner et al. (2022)). We have also shown that data-driven motility map models are highly predictive for multi-legged locomotion with or without slipping in both animals and robots as demonstrated by Zhao et al. (2022). Follow-up work by Bittner also showed that these models could be updated with online learning (Deng et al. (2024)).
However, data-driven motility map models remained either “easy” to identify near a periodic gait, or several related gaits (Bittner and Revzen (2021)), or “exponentially hard” to identify, requiring an amount of data exponential in the number of DoF for an exhaustive model (Dai et al. (2016)) throughout the phase space.
1.3 Contribution
@@
    We reformulate the problem of learning a motility map $A(r,\dot{r})$ as a multi-valued manifold learning problem, allowing for non-linearities in $\dot{r}$ such as those predicted by Kvalheim et al. (2019), other non-linear structures in $r$ and $\dot{r}$, and hysteretic multi-valued motility maps to be represented.
@@
    We provide a data transformation that allows a degree of linear extrapolation in the $\dot{r}$ direction to be controlled, capturing the fact that the manifold representing the motility map is often close to being a ruled surface with $A(r,\dot{r})=A(r)\cdot\dot{r}$.
@@
    We present an adaptation of well-known Gaussian Mixture Regression techniques to the machine learning problem of regressing a multi-valued function with a strong prior towards linearity in some parameters.
@@
    We demonstrate that when applied to the datasets of Zhao and Revzen (2020) and Zhao (2021) (available at BIRDS-Lab (2021a), BIRDS-Lab (2021b) respectively), consisting of multi-legged robot data, this new approach noticeably improves the predictive ability relative to the prior state of the art.
2 Methods
2.1 Manifold Learning
Manifold learning is the process of estimating the structure of a manifold from a set of noisy observations of points taken from that manifold, usually while the manifold is embedded in a larger space and the noise can offset the observations in that larger space. This form of learning operates on the assumption that high-dimensional data often lies on or near a lower-dimensional manifold embedded within the higher-dimensional space. Since every smooth function defines and is uniquely described by a manifold which is its graph, methods of estimating manifolds applied to the input-output pairs of a function are closely related to methods of function regression.
@@ 2.1 Manifold Learning
From a practical standpoint, many functions modeling physical systems are applied to a continuous input – either time, or some continuous function of time, and can be multi-valued when treated as a function of the current input. Such multi-valued functions take on different values for the same input depending on the path used to reach it. The most common example of these functions are hystereses, such as blacklash in gear mechanisms, diode hysteresis in electronics, and robot contact force hysteresis in contacts made through rubber-padded limbs. To capture this kind of multi-valued behavior one may build a model $y=f(x,y_{p})$ where the output $y$ is taken as a function of $x$ and a recent output $y_{p}$, usually arising from the previous time-step [^3] This can also be thought of as a dynamical model $y_{n+1}=f(y_{n};x)$ where $x$ is a control input, constructed over dynamics where the control is “very high gain”, allowing for only a few values of $y$ to be possible for any control $x$. We do not further explore this perspective here.. Using this approach, the model “chooses” a branch of the possible multi-valued outputs and remains on that branch. If an observer estimating the true value of $y$ is used to inform the value of $y_{p}$ and $(x,y_{p})$ is near a branching point where $y$ can bifurcate as $x$ is driven through it, the model will follow whichever branch of values is most consistent with the observed $y$ — a highly useful property in practice. Note that if we require the condition $f(x,y_{p})\to y$ for all $y_{p}\to y$, the function $f$ will be continuous; this will come into play below.
Here we suggest that the motility map’s graph be considered as such a multi-valued function, and we attempt to model it using Gaussian Mixture Regression (GMR).
2.2 Gaussian Mixture Regression of Motility Maps
In this section, we mathematically describe how a Gaussian Mixture Model (GMM-s) and the closely related technique of Gaussian Mixture Regression (GMR) are able to model a motility map.
Consider the following smooth manifold representing the graph of the motility map $A(r)$ in an ambient space $X$:
$P$ =$=\{(v_{b},r,\dot{r})\in X|v_{b}=A(r)\dot{r}\}$ $X:=\mathbb{R}^{N_{b}+2N_{s}},$ (2)
where $N_{b},N_{s}$ are the dimension of body space and shape space respectively. We assumed there exists some probability measure $d\mu_{P}(p)$ on $P$ representing the sampling process that generated our data in $P$, and some “measurement error” probability distribution $\text{Pr}(x|p)$ on the ambient space for each point $p\in P$. This defines a measurement probability distribution on the ambient space
[…]
Our approach is to approximate the density $\Pr$ with a Gaussian Mixture Model (McLachlan and Peel (2000)), and use the predictions obtained from these Gaussian components to provide a Gaussian Mixture Regression value for $A(r)\dot{r}$. Following Sung (2004) we note that were we only predicting outputs from inputs (without additional recent inputs), this would be a direct application of the Nadaraya-Watson estimator (Nadaraya (1964); Watson (1964)), and it is established that Nadaraya-Wastson estimators are universal estimators for continuous functions .
2.2.1 Gaussian Branching Regression
For completeness, we describe our approach from first principles, building up from the individual Gaussian components of the GMM modeling the graph.
@@ 2.2.1 Gaussian Branching Regression
Take a linear model $y=Ax$, $x\in X$, and $y\in Y$, applied to a sample of $x$ points coming from a Gaussian distribution $N(\mu,\sigma)$. The graph of this linear model is a point set in the direct sum vector space $W:=X\oplus Y$. We use the direct sum operator $\oplus$ to remind the reader that the vector space structure of the spaces $X$ and $Y$ extends naturally to $W$ as $(x\oplus y)+\alpha(v\oplus u):=(x+\alpha v)\oplus(y+\alpha u)$. Additionally, any linear map $M$ from $W$ to itself, i.e. $M\in\text{Aut}(W)$ has four parts: $M_{xx}\in\text{Aut}(Y)$, $M_{yy}\in\text{Aut}(Y)$, $M_{xy}\in L(Y,X)$ and $M_{yx}\in L(X,Y)$, such that $M(x\oplus y):=(M_{xx}x+M_{xy}y)\oplus(M_{yx}x+M_{yy}y)$ following the usual block structure of matrix operations. When $M_{xy}$ and $M_{yx}$ are both zero, we write the block diagonal $M$ as $M_{xx}\oplus M_{yy}$, noting that in this case $(M_{xx}\oplus M_{yy})(x\oplus y)=(M_{xx}x)\oplus(M_{yy}y)$.
@@
Given the distribution of $x$ is Gaussian, the graph of points $(x,Ax)$ is itself a sample from a (degenerate) Gaussian $N(\mu\oplus A\mu,S)$, where $S$ is a rank $\dim X$ symmetric non-negative matrix whose $S_{xx}$ block is $\sigma$ and whose $S_{yy}$ block is $A\sigma A^{\mathsf{T}}$. Assuming the measurements of $x$ (resp. $y$) are corrupted by (independent) Gaussian measurement noise with covariance $C_{x}$ (resp. $C_{y}$) the noisy data of the graph is still a sample from a Gaussian, now $N(\mu\oplus A\mu,~{}S+(C_{x}\oplus C_{y}))$, and no longer degenerate.
@@
Assuming $\sigma$ is sufficiently larger than $C_{x}$ and $C_{y}$, the matrix $A$ can be recovered from $S$ by using Total Least Squares regression. The TLS computation splits $W$ space along the eigenspaces of $S$, by decomposing $S=U^{\mathsf{T}}dU$ with $U$ orthogonal and $d$ a non-negative diagonal matrix with non-increasing elements along the diagonal. The TLS estimate of the graph is the kernel of the $\dim Y$ last rows of $U$, i.e. the space $U_{yx}x+U_{yy}y=0$, or
@@
Given a Gaussian input distribution and Gaussian measurement noise for input and output we conclude that the measurements of the resulting graph follow a Gaussian distribution as well. Furthermore, with sufficiently small measurement noise this graph Gaussian can be used to estimate the output for any input by treating the graph Gaussian’s covariance as a linear constraint.
To estimate a multi-valued function we combined the estimates provided by different components of the GMM of its graph based on their respective weights. For classical GMR, the weights would be $w_{i}\Pr(x|N(\mu_{i},\sigma_{i}))$ – the weight $w_{i}$ of the Gaussian component in the mixture, multiplied by the probability of the $x$ value for that Gaussian component. Instead we used $w_{i}\Pr((x,y_{p})|N(\mu_{i},\sigma_{i}))$, taking into account the previous output $y_{p}$. When multiple branches of $y$ values existed for an $x$, and were distinct with respect to the measurement noise, the Gaussians corresponding to $y$ values on branches other than that of $y_{p}$ had vanishingly small probabilities and thus little to no influence on the interpolation results. The result is what we call a “Gaussian Branching Regressor (GBR)” which is defined on input of the form $(x,y_{p})$ instead of $x$, and sensibly selects a branch and remains within it when computing a path integral of e.g. the motility map from data that seems to be multi-valued.
2.2.2 An illustrative example
@@
Consider the algebraic variety $y^{2}=(x-1)^{2}(x+1)^{2}$ shown in Fig. 1. It consists of two branches: $y=(x-1)(x+1)$ and $y=-(x-1)(x+1)$, each of which is a manifold, and provides us with a toy example of a multi-valued function. The variety itself as a point set in the $(x,y)$ plane is of measure 0, but a sampling process sampling $x$ uniformly in $[-1.5,1.5]$, then taking a corresponding $y$ from either branch at equal probability, and finally corrupting both $x$ and $y$ with $N(0,0.15)$ Gaussian measurement noise, produces a measurable probility distribution on the $(x,y)$ plane. We built a GMM model of these data using sklearn.mixture.GaussianMixture, which produced a collection of Gaussian components $N(\mu_{k},S_{k})$ with their respective weights $w_{k}$. We applied equation 4 for each $S_{k}$ to obtain the slope $A_{k}$. Note that $\mu_{k}$ is 2D, i.e. $\mu_{k}:=(\mu_{x,k},\mu_{y,k})$.
Our GBR predicted value $\hat{y}(x,y_{p})$ for the function at $x$ given the previous value of $y_{p}$ is:
$c_{k}$ $:=w_{k}\Pr((x,y_{p})|N(\mu_{k},S_{k}))$ (5)
$\hat{y}(x,y_{p})$ $:=\frac{1}{\sum_{k}c_{k}}\sum_{k}c_{k}\left(A_{k}(x-\mu_{x,k})+\mu_{y,k}\right)$ (6)
@@
Figure 1: Example of Gaussian Branching Regression (GBR). We considered the algebraic variety $y^{2}=(x-1)^{2}(x+1)^{2}$, which has symmetric positive (blue line) and negative (red line) branches. We simulated 1000 noisy measurements (blue circles) with measurement noise $N(0,0.15)$ in both $x$ and $y$ with $x$ taken uniformly from $[-1.5,1.5]$, and the $y$ branch chosen at random with equal probability. Using scikit.learn.gmm, we represented the sample points as a GMM and plotted the covariance ellipses of the constituent Gaussians (black ellipses). We then computed and plotted the predicted value of the model for a point traveling from $(-1,1)$ to the right, thus taking the positive (purple) branch, and similarly for a point traveling to the right from $(-1,-1)$ on the negative (dark red) branch.
2.3 Pre-processing to Produce a Ruled Surface
As noted above, GBR is quite general. It includes the ability to handle multi-valued functions, and because it is merely the application of the Nadaraya-Watson regressor to $f(x,y_{p})$ it inherits the universality of this regressor. The motility map, on the other hand, has some important additional structure that may be worth enforcing: its graph is a ruled surface since $v_{b}\ A(r)\dot{r}$ is linear in $\dot{r}$.